OpenStudy (anonymous):

The mean score for a standardized test is 1700 points. The results are normally distributed with a standard deviation of 75 points. If 10,000 students take the exam, how many would you expect to score above 1850 points?

OpenStudy (e.cociuba):

(1850 - 1700) / 75 = 150 / 75 = 2... 2 on the z-table equals to 0.9772. But this is for 1850 and below, we want above so P[z > 2] = 1 - P[z < 2] = 1 - 0.9772 = 0.0228 10000 * 0.0228 = 228, approximately 228 people will score above 1850 ANSWER: 228 :)